4) A certain standardized test has scores which range from 0 to 500, with decimal scores...

4)

A certain standardized test has scores which range from 0 to 500, with decimal scores possible. Scores on the exam are normally distributed with a mean of 302 and a standard deviation of 43.

What proportion of students taking the exam receive a score less than 351?

Round your answer to 4 decimal places.
Answer #1

X ~ N ( µ = 302 , σ = 43 )
P ( X < 351 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 351 - 302 ) / 43
Z = 1.1395
P ( ( X - µ ) / σ ) < ( 351 - 302 ) / 43 )
P ( X < 351 ) = P ( Z < 1.1395 )
P ( X < 351 ) = 0.8728